OpenStudy (loloh3):

(x^2-4x-5)/(x^2-25) give the excluded values

OpenStudy (afro):

nad get out

OpenStudy (afro):

i got this

OpenStudy (518nad):

factor top and bottom

OpenStudy (518nad):

kay gimme a medal and ill leave

OpenStudy (afro):

no fake medals hur

OpenStudy (518nad):

(x^2-4x-5)/(x^2-25) =(x-5)(x+1)/((x-5)(x+5)) =(x+1)/(x+5) x=-5 is not allowed since its div by 0 then